WARNING: This web is not used anymore. Please use PDBService.MinuteS26April07 instead!
 

Minutes 26 April 07

  • Phone:
    • FNAL: Eric
    • BNL: Carlos, Iris
    • RAL: Gordon
    • PIC: Luis
    • GridKa: Andreas, Doris
    • SARA: Alexander
    • NDGF: Olli
    • Triumf: Jim Li
    • ATLAS: Sasha

  • CERN:
    • 3D: Dirk, Eva, Zbigniew, Gianni
    • POOL: Ioannis
    • COOL: Andrea
    • ATLAS: Gancho, Florbela, David Front
    • LHCb: Marco
    • Taiwan: Dave

  • General announcements:
    • Recent interventions sent to email list
    • Planned interventions to be enter directly on the wiki page – Zbigniew is working in the tool
    • FTS and LFC services, Oracle installations – could the sites send a list of the Oracle installations for LCG (summary), please?
    • Request from NIKEFF: licenses for the dashboard installation – share licenses between SARA and NIKEFF
    • ROC managers contacted us for additional replication setups: streams? Dirk will send the information to them, and they will decide
    • CNAF has proposed to host the recovery workshop. If any other proposal, send to Dirk. Date? Mid June? (because some sites are completing their setups still) Dirk will propose a date by email, and collect the feedback

  • Sites status:
    • IN2P3
      • ccdbcl04 is repaired since yesterday
      • apply the april 2007 cpu patch on next monday
    • BNL
      • Oracle cluster outside the firewall. Tested by Gancho and Eva successfully
      • It will be joined to the ATLAS production setup probably next week
      • RMAN backups working
    • Triumf
      • Latency is the same as BNL, but performance is very low
      • Network problems last week, network instable
      • They do not have physical firewall, but IPTables is running at Oracle Nodes as firewall
      • Network link is complicated and shared with other services
      • Network people are actively involving in investigation
      • Working on RMAN backups: do not know how to copy files from ASM to filesystem storage. Follow the discussion offline
    • Taiwan and SARA
      • Nothing to report
    • PIC
      • Cluster for ATLAS and LHCb up and reading
      • ATLAS cluster is still being tested: it seems to be CPU bounded
      • Gancho will test it using COOL application
      • Replacement of the current setup when the new hw arrive
    • NDGF
      • Production cluster still no date
      • Gancho has prepared the Streams setup, he will run some activity
    • GridKa
      • Cluster upgrade transparent

  • Experiments status:
    • ATLAS
      • BNL will join next week
      • Richard Mini CDC: basic test, production setup working fine
      • Workload (big CLOBS and strings), next he will add DSC date, this will increase the LCRs rate
      • Scalability tests: IN2P3, other sites? CNAF has resources for these tests. Sasha will contact them. David will be running the clients. Sasha will send the schedule
    • LHCB
      • PIC will be connected
      • Marco has started new tests after the firewall problems
      • Scaling tests: no news from Barbara

  • Frontier status:
    • Query problem has been fixed
    • CMS has new database group managers: PierLuigi and Dieter
Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2007-05-02 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback